OpenStudy (anonymous):

given that y = x^(1/3), use calculus to determine an approximate value for 8030^(1/3). pls show working.

OpenStudy (dumbcow):

why would you need calculus? cubed root of 8030 is approx 20.025

OpenStudy (hunus):

Using linear approximation \[L(a)=f(a)+f'(a)(x-a)\] Taking a to be 8000 (because we know the cube root of 8000 and it's close to 8030) \[y=f(x)=x^{1/3}\] \[f'(x)=1/(3x^{2/3})\] \[L(8000)=f(8000)+f'(8000)(x-8000)\] \[L(8000)=(8000)^{1/3}+1/(3(8000)^{2/3})(x-8000)\] \[L(8000)=20+(1/1200)(x-8000)\] \[L(8000)=20 +x/1200-20/3\] Thus the linear approximation of y=x^1/3 near 8000 is \[L(x)=40/3+x/1200\] \[L(8030)=40/3+8030/1200\] Which is 20.025

OpenStudy (anonymous):

thats would be 2.0

OpenStudy (anonymous):

you dumb cow, the point of the question is to show we know calculus. not to prove that we know how to use a calculator.